OpenStudy (anonymous):

integral ((x^2 + 1)^1/2)/x from 1 to 2

OpenStudy (anonymous):

u=x^2+1 du =2x dx du/2x=dx

OpenStudy (anonymous):

this is really crazy looking but can't you u just be be the top

OpenStudy (anonymous):

I would use u substitution here. Try making: u=(x^2+1)

OpenStudy (anonymous):

\[\int\limits_{1}^{2} \sqrt{x ^{2} - 1} / x\]

OpenStudy (anonymous):

I'm supposed to do it by setting up a right triangle of some sort

OpenStudy (anonymous):

x=(u-1)^(1/2)

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

trignometric substitution

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

u= asec(theta)

OpenStudy (anonymous):

let x = 1sec(theta)

OpenStudy (anonymous):

dx=sec(x)tan(x)

OpenStudy (anonymous):

how did you come up with =1sec(theta?

OpenStudy (anonymous):

there is special instances ... like how i remember is that when you were taught inverse trig integrals

OpenStudy (anonymous):

i'm not sure how they work again but it is triangles that describe the relationship

OpenStudy (anonymous):

for inverse A^2+u^2 =was how you noticed arctan correct well the same for trig substitution only with a squareroot

OpenStudy (anonymous):

when you did inverse trig when you saw sqrt of u^2-a^2.... secant was the first you thought of

OpenStudy (anonymous):

http://www.youtube.com/watch?v=3lC5AuCFK4c - here is a video on it

OpenStudy (anonymous):

oh okay, so in this problem, x would be hypotenuse, adjacent would be 1, and opposite would be root x^2 -1?

OpenStudy (anonymous):

yes once you get all the way at the end you have to remember you used substitution so you have to put it back into x form

OpenStudy (anonymous):

okay so sec (theta) = x, so dx = sec(theta)tan(theta)

OpenStudy (anonymous):

tan (theta) = root x^2 - 1

OpenStudy (anonymous):

then substituting: tan(theta)sec(theta)tan(theta)/sec(theta)

OpenStudy (anonymous):

so far looks good

OpenStudy (anonymous):

so cancel sec(theta), and then its just the integral of tan^2 (theta)? is that right?

OpenStudy (anonymous):

i believe so

OpenStudy (anonymous):

how do i do integral of tan^2 theta?

OpenStudy (anonymous):

however you cannot take the integral of tan^2 so you are going to have to do

OpenStudy (anonymous):

trignometric integrals which is what you were doing before

OpenStudy (anonymous):

okay how would i start that?

OpenStudy (anonymous):

with the identity \[\cos^2(x)+\sin^2(x)=1\] \[1+\tan^2(x)=\sec^2(x)\]

OpenStudy (anonymous):

okay so tan^2 x = sec^2 x -1

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

so the integral of that is tan x - x?

OpenStudy (anonymous):

well now you integrated it... but however you used substituion much like you did with u..... so when you evaluate you have to do something

OpenStudy (anonymous):

and your x should be thetas

OpenStudy (anonymous):

oh yeah thanks

OpenStudy (anonymous):

\[\tan \theta - \theta from 1 \to 2\]

OpenStudy (anonymous):

so would i substitute something in for theta?

OpenStudy (anonymous):

those are the integration limits for x

OpenStudy (anonymous):

you can either change your integration limits to theta or x

OpenStudy (anonymous):

ill change theta to x, its easier for me

OpenStudy (anonymous):

what would theta change to though?

OpenStudy (anonymous):

arcsec

OpenStudy (anonymous):

use your substitution that u started with to solve for theta

OpenStudy (anonymous):

1.22

OpenStudy (anonymous):

okay so: \[(\sqrt{x^{2}} - 1) - arcsec x from 1 t\]

OpenStudy (anonymous):

that looks correct

OpenStudy (anonymous):

okay how do i solve arcsec x now?

OpenStudy (anonymous):

integral sqrt(1+x^2)/x dx For the integrand, sqrt(x^2+1)/x substitute x = tan(u) and dx = sec^2(u) du. Then sqrt(x^2+1) = sqrt(tan^2(u)+1) = sec(u) and u = tan^(-1)(x): = integral csc(u) sec^2(u) du For the integrand csc(u) sec^2(u), use the trigonometric identity sec^2(u) = tan^2(u)+1: = integral (tan^2(u)+1) csc(u) du Expanding the integrand (tan^2(u)+1) csc(u) gives csc(u)+tan(u) sec(u): = integral (csc(u)+tan(u) sec(u)) du Integrate the sum term by term: = integral csc(u) du+ integral tan(u) sec(u) du For the integrand tan(u) sec(u), substitute s = sec(u) and ds = tan(u) sec(u) du: = integral 1 ds+ integral csc(u) du The integral of csc(u) is -log(cot(u)+csc(u)): = integral 1 ds-log(cot(u)+csc(u)) The integral of 1 is s: = s-log(cot(u)+csc(u))+constant Substitute back for s = sec(u): = sec(u)-log(cot(u)+csc(u))+constant Substitute back for u = tan^(-1)(x): = sqrt(x^2+1)-log((sqrt(x^2+1)+1)/x)+constant An alternative form of the integral is: = sqrt(x^2+1)-csch^(-1)(x)+constant Which is equivalent for restricted x values to: = sqrt(x^2+1)-log(sqrt(x^2+1)+1)+log(x)+constant now you aply the limits u'll get the answer as 1.22

OpenStudy (anonymous):

the book said by using trig substitution...

OpenStudy (anonymous):

s this ur ans!!!!1

OpenStudy (anonymous):

okay im kinda stuck, I'm at :\[\sqrt{3 - arcsec 2} - \sqrt{arcsec 1}\]

OpenStudy (anonymous):

shouldn't arc bet outside the sqr rt

OpenStudy (anonymous):

wait a second...

OpenStudy (anonymous):

okay so i tan theta - theta from 1 to 2 is what we are trying to solve

OpenStudy (anonymous):

arccos(1/x)

OpenStudy (anonymous):

yess from there you are correct

OpenStudy (anonymous):

and tan theta = root x^2 - 1 and theta is arc sec x

OpenStudy (anonymous):

so yes i messed up there

OpenStudy (anonymous):

\[\sqrt{3} - arcsec 2 - acrsec 1\] is the answer?

OpenStudy (anonymous):

i believeso let met check with my imaginary calculator... let me go get it

OpenStudy (anonymous):

it doesn't come out right when i do it

OpenStudy (anonymous):

i know tan theta - theta is correct

OpenStudy (anonymous):

yeah same, I'm just not sure after that

OpenStudy (anonymous):

you could always go into theta

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

x=sec(x) correct?

OpenStudy (anonymous):

so 2 = sec(theta) and 1=sec(theta)

OpenStudy (anonymous):

sec(theta) = x

OpenStudy (anonymous):

will get you the limits in theta

OpenStudy (anonymous):

find the angles in which sec(theta) =1 and 2

OpenStudy (anonymous):

one will be when cos=1/2

OpenStudy (anonymous):

other will be when cos =1

OpenStudy (anonymous):

so 0 and pi/3

OpenStudy (anonymous):

are your upper and lower

OpenStudy (anonymous):

okay so then what is the equation?

OpenStudy (anonymous):

if you switch the limits you don't have to take it out of theta

OpenStudy (anonymous):

so tan(theta)-theta)]from 0 to pi/3

OpenStudy (anonymous):

did your teacher ever teach you that? you can keep your u substitutions in u form if you plug x into your substitution

OpenStudy (anonymous):

so root 3 - pi/3 is the answer?

OpenStudy (anonymous):

i think that is right actually!

OpenStudy (anonymous):

yes that looks right idk how she got 1.22

OpenStudy (anonymous):

yeS!!!! you are the best

OpenStudy (anonymous):

glad i was help

OpenStudy (anonymous):

btw i believe the way i solved your other questions is the way your teacher wants you to solve it as in my book, trig integrals is right before trig substitution

OpenStudy (anonymous):

and parts you probably don't even have a clue yet

OpenStudy (anonymous):

okay thank you!